OpenStudy (anonymous):

Solve: 2(x + 8) − (x + 1) = 20

OpenStudy (anonymous):

What would you do first?

OpenStudy (saifoo.khan):

2x + 16 -x -1 = 20

OpenStudy (anonymous):

2x+16 -x - 1 = 20 x + 15 =20 x = 5

OpenStudy (anonymous):

2x+16-x-1=20 x=5